# Multiscale Hodge Scattering Networks for Data Analysis

### v5

@@ 2.3 The $\kappa$-HGLET
The $\kappa$-HGLET is a generalization of the Hierarchical Graph Laplacian Eigen Transform (HGLET) [30] from functions on the nodes of a graph, to functions on the $\kappa$-simplices in a given simplicial complex [7]. The HGLET, in turn, can be viewed as a generalization of the Hierarchical Block Discrete Cosine Transform (HBDCT), which is generated by creating a hierarchical bipartition of the signal domain and computing the DCT of the local signal supported on each subdomain.
Let $\{{\boldsymbol{\phi}}^{j}_{k,l}\}$ be the set of basis vectors in the $\kappa$-HGLET dictionary where $j$ denotes the level of the partition (with $j=0$ being the root), $k$ indicates the partition within the level, and $l$ indexes the elements within each partition in increasing frequency. Let $C^{j}_{k}$ refer to the $\kappa$-region consisting of the support of partition $k$ at level (or scale) $j$, and let $n^{j}_{k}\,:=\,\left|C^{j}_{k}\right|$. Hence $C^{0}_{0}=C_{\kappa}$ and $n^{0}_{0}=|C_{\kappa}|\,=:\,n$.
In order to compute the transform, we first compute the complete set of eigenvectors $\{{\boldsymbol{\phi}}^{0}_{0,l}\}_{l=0:n^{0}_{0}-1}$ of $\mathcal{L}=\mathcal{L}(C_{0}^{0})$, and order them by nondecreasing eigenvalues. We then partition $C^{0}_{0}$ into two disjoint $\kappa$-regions $C^{1}_{0}$ and $C^{1}_{1}$ by forming the Fiedler vector of $\mathcal{L}$. We note that: 1) one can use any other bipartition methods; and 2) bipartitioning with the Fiedler vector in the $\kappa$-region setting requires additional steps vs. the graph setting, because of its rather intricate behaviors; see [7] for the details. We iterate the same procedure for $C^{1}_{0}$ and $C^{1}_{1}$ to generate the eigenvectors $\{{\boldsymbol{\phi}}^{1}_{0,l}\}_{l=0:n^{1}_{0}-1}$ and $\{{\boldsymbol{\phi}}^{1}_{1,l}\}_{l=0:n^{1}_{1}-1}$. Note that $n^{1}_{0}+n^{1}_{1}=n^{0}_{0}=n$, and that all of the elements in $\{{\boldsymbol{\phi}}^{1}_{0,l}\}$ are orthogonal to those in $\{{\boldsymbol{\phi}}^{1}_{1,l}\}$ since their supports are disjoint. The set $\{{\boldsymbol{\phi}}^{1}_{0,l}\}_{l=0:n^{1}_{0}-1}\cup\{{\boldsymbol{\phi}}^{1}_{1,l}\}_{l=0:n^{1}_{1}-1}$ form an ONB for vectors in $X_{\kappa}$. From here, we apply this process recursively, generating an ONB for each level in the given hierarchical bipartition tree. If the hierarchical bipartition tree terminates at every region containing only a $\kappa$-simplex singleton, then the final level will simply be the standard basis of ${\mathbb{R}}^{n}$. Each level of the dictionary contains an ONB whose vectors have the support of roughly half the size of the previous level. There are roughly $(1.5)^{n}$ possible ONBs formed by selecting different covering sets of regions from the hierarchical bipartition tree. We also note that the computational cost of generating the entire dictionary is $\mathcal{O}(n^{3})$. See [7] for the actual algorithm to generate the $\kappa$-HGLET on a given $C_{\kappa}$ and further details.
@@ 2.4 The $\kappa$-GHWT
This basis dictionary is based on the Generalized Haar-Walsh Transform (GHWT) [8], which can itself be viewed as a generalization of the Haar-Walsh wavelet packets [31, Sec. 8.1]. This is formed by first generating a hierarchical bipartition tree of $C_{\kappa}$ as for the $\kappa$-HGLET. We then work in a bottom-up manner, beginning with the finest level $j={j_{\mathrm{max}}}$ where each region only contains a single element that is the indicator vector of that region. We call them scaling vectors and label them as $\{{\boldsymbol{\psi}}^{{j_{\mathrm{max}}}}_{k,0}\}_{k=0:n-1}$. For the next level $j={j_{\mathrm{max}}}-1$, we first assign a constant scaling vector for the support on each region. Then, for each region that contains two children in the partition tree, we form a Haar vector by subtracting the scaling vector of the child element with a higher index from that of the child element with a lower index. This procedure will form an ONB $\{{\boldsymbol{\psi}}^{{j_{\mathrm{max}}}-1}_{k,l}\}_{k=0:K-1,l=0:l(k)-1}$ (where $K$ is the number of $\kappa$-regions at level ${j_{\mathrm{max}}}-1$ and $l(k)=1$ or $2$ depending on the region $k$) whose vectors have support of at most 2. For the level $j={j_{\mathrm{max}}}-2$, we begin by computing the scaling and Haar vectors as before. Next, for any region that contains three or more elements, we also compute Walsh vectors by adding and subtracting the Haar vectors of the children regions. From here, we form the rest of the dictionary recursively. A full description of this algorithm is given in [30] for the $\kappa=0$ case and in [7] for the general case of $\kappa>0$.
Note that like the $\kappa$-HGLET, each level of the dictionary forms an ONB, and at each level, basis vectors have the support of roughly half the size of the parent level. These basis vectors also have the same support as the corresponding $\kappa$-HGLET basis vectors (that is, $\operatorname{supp}({\boldsymbol{\phi}}^{j}_{k,l})=\operatorname{supp}({\boldsymbol{\psi}}^{j}_{k,l})$ for all $j,k,l$). However, the computational cost of computing the $\kappa$-GHWT is only $\mathcal{O}(n\log n)$.
@@ 3 Multiscale Hodge Scattering Transform
Let the $\kappa$-HGLET or $\kappa$-GHWT dictionary vectors be arranged as ${\boldsymbol{\Phi}}^{J}\,:=\,\left\{\Phi^{j}\right\}_{j=0}^{J}$ where each $\Phi^{j}$ is an ONB at scale (or level) $j$ with $j=0$ being the finest scale basis, composed of delta functions. Note that this definition of the scale parameter $j$ is the opposite of that used in the previous sections. In general, we have ${j_{\mathrm{max}}}+1\approx\log_{2}n+1$ different levels given by the hierarchical bipartition scheme, but in practice, the features extracted by large $j$ values are not very descriptive [4]. Hence, we typically use the first $J+1(\leq{j_{\mathrm{max}}}+1)$ levels.
@@
Let ${\boldsymbol{f}}\in X_{\kappa}$ and let $[{\boldsymbol{f}}]_{i}$ denote ${\boldsymbol{f}}$ evaluated at simplex $i$. In addition, let us write $\left|{\boldsymbol{f}}\right|^{q}\,:=\,\left(\left|[{\boldsymbol{f}}]_{1}\right|^{q},\ldots,\left|[{\boldsymbol{f}}]_{n}\right|^{q}\right)\in{\mathbb{R}}^{n}$ for $q\in{\mathbb{N}}$. We propose to compute the $q$th moment up to some maximum $Q\in{\mathbb{N}}$ of the 0th-order and 1st-order scattering coefficients:
[…]
and the 2nd-order scattering coefficients:
[…]
And higher-order moments can be computed in a similar manner:
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right){\boldsymbol{f}}\,:=\,\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j^{(m)}}\left|\Phi^{j^{(m-1)}}\left|\cdots\left|\Phi^{j^{(1)}}{\boldsymbol{f}}\right|\cdots\right|\right|\right|^{q}\right]_{i},$ (5)
where $0\leq j^{(1)}<\cdots<j^{(m)}\leq J$. However, due to the combinatorial blow-up in the number of features at each order, it is rare to use more than 2nd or 3rd-order features. Note that the $m$th order features are computed by applying a multiplication with the appropriate (sparse) weight $\Phi^{j}$ then applying a pointwise nonlinearity to $(m-1)^{\text{st}}$-order features. Further, as we later find in our numerical experiments, high-order moments $(Q>4)$ are not very useful in practice due to their instability [2, 4, 25]. {hide} We notate the transform without the sum or normalization factor as $\hat{S}$. We can write the higher order features as:
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right){\boldsymbol{f}}$ $=S^{1}\left(q,j^{(m)}\right)\hat{S}^{m-1}\left(1,j^{(1)},\ldots,j^{(m-1)}\right){\boldsymbol{f}}=$ (6)
$\cdots=S^{1}\left(q,j^{(m)}\right)\hat{S}^{1}\left(1,j^{(m-1)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right){\boldsymbol{f}}$
By defining the operator $\hat{S}^{1}(j,q){\boldsymbol{f}}\,:=\,\left|\Phi^{j}{\boldsymbol{f}}\right|^{q}$, which maps ${\boldsymbol{f}}\in{\mathbb{R}}^{n}$ to ${\mathbb{R}}_{\geq 0}^{n}$, we can also rewrite the higher-order features Eq. (5) in a more comprehensible manner as:
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right){\boldsymbol{f}}=S^{1}\left(q,j^{(m)}\right)\hat{S}^{1}\left(1,j^{(m-1)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right){\boldsymbol{f}}$ (7)
@@
This behavior mimics the architecture of convolutional neural networks (with fixed weights) and has been studied extensively [1]. However, unlike traditional feed-forward networks, rather than only considering the features extracted at end of the cascade of linear and nonlinear layers, we concatenate the representation extracted at each layer for downstream analysis. In general, we refer to the process of extracting these features as the Multiscale Hodge Scattering Transform (MHST). Later we will analyze these features with relatively simple machine learning models such as support vector machines (SVMs) and logistic regression models (LRMs). When working with such models with learnable parameters in concert with these deterministic features, we refer to it as the Multiscale Hodge Scattering Network (MHSN).
@@ 3.1 Local Pooling
In general, we can gather all of the moments $\leq Q$ and of orders $\leq M$ to have a total of $Q\sum_{k=0}^{M}\binom{J+1}{k}$ features for a given signal. The summations from $i=1$ to $i=n$ in (3)–(5) can be viewed as global pooling operations. In situations where permutation invariance is not required (i.e., all signals are defined on a fixed complex with known node ordering), however, we can omit these sums, i.e., no pooling is done. As a result, we are left with $Qn\sum_{k=0}^{M}\binom{J+1}{k}$ features for each signal. We can also generate intermediate features between these two extremes: retain sums over each region $k$ at level $j$ instead of not summing at all or summing all the regions of level $j$ in (3)–(5). This can be viewed as local pooling operations and results in a tuple of features rather than a single value as in the original scattering transform. This is similar to the windowed scattering transforms recently proposed by [25], but here we leverage the multiscale decomposition provided by the hierarchical partition tree to avoid introducing a new user parameter to determine the window size. In the case of these local pools, we replace the normalization factor in the averaging operation with the number of elements in the local pool rather than the number in the entire simplex as defined in Eq. (8). This gives us a total of $Q\sum_{j=0}^{M}K^{j}\binom{J+1}{j}$ features, where $K^{j}$ is the number of local sums taken in the $j$th level, i.e., the number of regions at scale $j$.
We denote these transforms as $S^{m}_{j}$ where $j$ denotes the level at which the sum has taken place. So $S^{m}_{j_{\mathrm{max}}}=S^{m}$ indicated the standard max-pooling-like scheme as defined by [2, 4]. Then $S^{m}_{0}$ denotes the transform without any sums while $S^{m}_{2}$ denotes the transform with local pools determined by the third level (i.e., $j=2$) of the partition tree. In general we have:
$S^{m}_{j}\left(q,j^{(1)},\ldots,j^{(m)}\right){\boldsymbol{f}}\,:=\,\Biggl{\{}\frac{1}{n^{j}_{k}}\sum_{i\in C^{j}_{k}}\left[\left|\Phi^{j^{(m)}}\left|\Phi^{j^{(m-1)}}\left|\cdots\left|\Phi^{j^{(1)}}{\boldsymbol{f}}\right|\cdots\right|\right|\right|^{q}\right]_{i}\Biggr{\}}_{k=0}^{K^{j}-1}$ (8)
Note that the subscript $j$ indicates where the final averaging has taken place whereas $j^{(m)}$ indicates the index of the basis elements used to compute the feature.

### v6

@@ 2.3 The $\kappa$-HGLET
The $\kappa$-HGLET is a generalization of the Hierarchical Graph Laplacian Eigen Transform (HGLET) [32] from functions on the nodes of a graph to functions on the $\kappa$-simplices of a given simplicial complex [10]. The HGLET itself can be viewed as a generalization of the Hierarchical Block Discrete Cosine Transform (HBDCT), which is obtained by creating a hierarchical bipartition of the signal domain and computing the DCT of the local signal supported on each subdomain. For the reader’s convenience, we summarize the key points of the hierarchical bipartition tree of an input graph $G$ and the original HGLET for node signals in Appendices A and B, respectively.
Let $\{\bm{{\phi}}^{p}_{k,l}\}$ denote the set of basis vectors in the $\kappa$-HGLET dictionary, where $p$ is the level of the partition (with $p=0$ being the root), $k$ indexes the partition within the level, and $l$ indexes the elements within each partition in order of increasing frequency. Let $C^{p}_{k}$ be the $\kappa$-region corresponding to the support of partition $k$ at level $p$, and let $n^{p}_{k}\,:=\,|C^{p}_{k}|$. Thus $C^{0}_{0}=C_{\kappa}$ and $n^{0}_{0}=|C_{\kappa}|\,=:\,n$.
To compute the transform, we first obtain the complete set of eigenvectors $\{\bm{{\phi}}^{0}_{0,l}\}_{l=0:n^{0}_{0}-1}$ of $\mathcal{L}=\mathcal{L}(C^{0}_{0})$, ordered by nondecreasing eigenvalues. We then bipartition $C^{0}_{0}$ into two disjoint $\kappa$-regions $C^{1}_{0}$ and $C^{1}_{1}$ using the Fiedler vector of $\mathcal{L}$. We note that: 1) other bipartitioning methods may be used; and 2) bipartitioning with the Fiedler vector in the $\kappa$-region setting requires additional steps compared to the graph setting, due to its more intricate behavior (see [10] for details).
The same procedure is applied recursively to $C^{1}_{0}$ and $C^{1}_{1}$, yielding eigenvectors $\{\bm{{\phi}}^{1}_{0,l}\}_{l=0:n^{1}_{0}-1}$ and $\{\bm{{\phi}}^{1}_{1,l}\}_{l=0:n^{1}_{1}-1}$. Since $n^{1}_{0}+n^{1}_{1}=n^{0}_{0}=n$ and the supports of $\{\bm{{\phi}}^{1}_{0,l}\}$ and $\{\bm{{\phi}}^{1}_{1,l}\}$ are disjoint, these two sets are orthogonal. Their union $\{\bm{{\phi}}^{1}_{0,l}\}\cup\{\bm{{\phi}}^{1}_{1,l}\}$ forms an ONB for vectors in $X_{\kappa}$. Continuing this process down the hierarchical bipartition tree produces an ONB at each level. If the tree terminates with regions containing single $\kappa$-simplices, the final level coincides with the standard basis of ${\mathbb{R}}^{n}$.
Each level of the dictionary contains an ONB whose vectors have support roughly half the size of those at the previous level. There are approximately $(1.5)^{n}$ possible ONBs obtainable by selecting different covering sets of regions from the hierarchical bipartition tree. The computational cost of generating the entire dictionary is $\mathcal{O}(n^{3})$. See [10] for the complete algorithm to construct the $\kappa$-HGLET on a given $C_{\kappa}$ and for further implementation details.
@@ 2.4 The $\kappa$-GHWT
The $\kappa$-GHWT is a generalization of the Generalized Haar–Walsh Transform (GHWT) [11] from functions on the nodes of a graph to functions on the $\kappa$-simplices of a given simplicial complex [10]. The GHWT itself can be viewed as a generalization of the Haar–Walsh wavelet packets [33, Sec. 8.1]. As with the $\kappa$-HGLET, the construction begins by generating a hierarchical bipartition tree of $C_{\kappa}$.
The transform is computed in a bottom-up manner, starting from the finest level $p={p_{\mathrm{max}}}$, where each region contains a single $\kappa$-simplex represented by its indicator vector (i.e., one of the standard basis vectors of ${\mathbb{R}}^{n}$). These vectors are called scaling vectors and are denoted $\{\bm{{\psi}}^{{p_{\mathrm{max}}}}_{k,0}\}_{k=0:n-1}$. At the next level, $p={p_{\mathrm{max}}}-1$, we first assign a constant scaling vector supported on each region. For any region with two children in the partition tree, we form a Haar vector by subtracting the scaling vector of the higher-index child from that of the lower-index child. This yields an orthonormal basis (ONB) $\{\bm{{\psi}}^{{p_{\mathrm{max}}}-1}_{k,l}\}_{k=0:K-1,\,l=0:l(k)-1}$, where $K$ is the number of $\kappa$-regions at level ${p_{\mathrm{max}}}-1$ and $l(k)\in\{1,2\}$, with each vector supported on at most two simplices.
At level $p={p_{\mathrm{max}}}-2$, we again compute scaling and Haar vectors as before. In addition, for any region containing three or more elements, we construct Walsh vectors by adding and subtracting the Haar vectors of its child regions. This process is then applied recursively up the tree, producing the full dictionary. A complete description of the algorithm is given in [11] for the $\kappa=0$ case and in [10] for the general case $\kappa>0$.
Like the $\kappa$-HGLET, each level of the $\kappa$-GHWT dictionary forms an ONB, and the basis vectors at each level have support roughly half the size of those at the previous level. Moreover, the $\kappa$-GHWT basis vectors share the same supports as the corresponding $\kappa$-HGLET basis vectors, i.e., $\operatorname{supp}(\bm{{\phi}}^{p}_{k,l})=\operatorname{supp}(\bm{{\psi}}^{p}_{k,l})$ for all $p,k,l$. However, the computational cost of constructing the $\kappa$-GHWT is only $\mathcal{O}(n\log n)$, in contrast to the $\mathcal{O}(n^{3})$ cost of the $\kappa$-HGLET.
@@ 3 Multiscale Hodge Scattering Transform
To discuss our version of the scattering transform on a given simplicial complex, it is convenient to define the scale parameter $j\,:=\,{p_{\mathrm{max}}}-p$, where $p$ is the partition level introduced in the previous section. Thus, $j=0$ corresponds to the finest partition $p={p_{\mathrm{max}}}$, where each partition is a singleton whose basis vector is one of the standard basis vectors of ${\mathbb{R}}^{n}$, and $j={j_{\mathrm{max}}}\,:=\,{p_{\mathrm{max}}}$ corresponds to the coarsest scale at the root level $p=0$ (i.e., no partition) of the input simplicial complex. Let the $\kappa$-HGLET or $\kappa$-GHWT dictionary vectors be arranged as $\bm{{\Phi}}^{J}\,:=\,\left\{\Phi^{j}\right\}_{j=0}^{J}$ where each $\Phi^{j}\in{\mathbb{R}}^{n\times n}$ is an ONB at scale $j$ whose rows are the basis vectors. In general, the hierarchical bipartition scheme yields ${j_{\mathrm{max}}}+1\approx\log_{2}n+1$ distinct levels, but in practice features extracted at large $j$ are often less descriptive [5]. Hence, we typically use the $J+1$ finest levels with $0\leq J\leq{j_{\mathrm{max}}}$.
Let $\bm{f}\in X_{\kappa}$ and write $\bm{f}]_{i}$ for its value on simplex $i$. For $q\in{\mathbb{N}}$, define
@@
$\left|\bm{f}\right|^{q}\,:=\,\left(\left|[\bm{f}]_{1}\right|^{q},\ldots,\left|[\bm{f}]_{n}\right|^{q}\right)^{\scriptscriptstyle{\mathsf{T}}}\in{\mathbb{R}}^{n}$. We compute the $q$th moments (up to some maximum $Q\in{\mathbb{N}}$) of the 0th- and and 11st-order scattering coefficients:
$S^{0}(q)\bm{f}\,:=\,\frac{1}{n}\sum_{i=1}^{n}[\bm{f}]_{i}^{q},\,\quad S^{1}(q,j)\bm{f}\,:=\,\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j}\bm{f}\right|^{q}\right]_{i},\,\,0\leq j\leq J;1\leq q\leq Q\,$, (1)
and the 22nd-order scattering coefficients:
$S^{2}\left(q,j,j^{\prime}\right)\bm{f}\,:=\,\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j^{\prime}}\left|\Phi^{j}\bm{f}\right|\right|^{q}\right]_{i},\,\,0\leq j<j^{\prime}\leq J,\,\,1\leq q\leq Q\,.$ (2)
Higher-order scattering coefficients are defined similarly:
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right)\bm{f}\,:=\,\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j^{(m)}}\left|\Phi^{j^{(m-1)}}\left|\cdots\left|\Phi^{j^{(1)}}\bm{f}\right|\cdots\right|\right|\right|^{q}\right]_{i},$ (3)
where $0\leq j^{(1)}<\cdots<j^{(m)}\leq J$. Because the number of features grow combinatorially with the order $m$, it is uncommon to use orders beyond $m=2$ or $3$. Moreover, as our numerical experiments indicate, higher moments ($q>4$) are typically less useful in practice due to instability [3, 5, 28].
For clarity, define the layer operator $\hat{S}^{1}(j,q)\bm{f}\,:=\,\left|\Phi^{j}\bm{f}\right|^{q}$, which maps $\bm{f}\in{\mathbb{R}}^{n}$ to ${\mathbb{R}}_{\geq 0}^{n}$, so that the $m$th-order feature in Eq. (3) can be written succinctly as
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right)\bm{f}=S^{0}(q)\hat{S}^{1}\left(1,j^{(m)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right)\bm{f}$ (4)
@@
This mirrors the architecture of a convolutional neural network with fixed weights, as studied in [1]. Unlike conventional feed-forward networks, in which only the input to the final layer (i.e., the classification head)—the features extracted at the end of the cascade of linear and nonlinear transformations—is used for classification, we concatenate the features from every layer of the cascade for downstream analysis. We refer to the process of extracting these features as the Multiscale Hodge Scattering Transform (MHST). When these deterministic features are used in conjunction with a separate learnable model, e.g., support vector machine (SVM) or logistic regression model (LRM), we call the system a Multiscale Hodge Scattering Network (MHSN).
Figure 2 illustrates our MHSN ($q=1$) with global pooling on a toy $2$-complex with nine elements. The leftmost column shows $2$-GHWT bases at successive scales, with colors indicating the sign of the basis vector components (yellow = positive; purple = negative; green = 0). For $m=0$, $\hat{S}^{0}\bm{f}\in{\mathbb{R}}^{9}_{\geq 0}$ is computed by taking inner products of the input signal with the finest scale ($j=0$) $\delta$-function basis; the resulting coefficients are then averaged over all simplices—this averaging over the whole domain is what we call global pooling—to produce a scalar $S^{0}\bm{f}\in{\mathbb{R}_{\geq 0}}$. For $m=1$, $\hat{S}^{1}(j)\bm{f}\in{\mathbb{R}}^{9}_{\geq 0}$ is obtained by taking the inner products of $\hat{S}^{0}\bm{f}$ (identical to $\bm{f}$) with the scale-$j$ GHWT basis vectors, then applying pointwise modulus (absolute value). Global pooling yields the scattering coefficient $S^{1}(j)\bm{f}\in{\mathbb{R}_{\geq 0}}$. Repeating this for each $j=0,1,\ldots,J$, produces $S^{1}\bm{f}\in{\mathbb{R}}^{J+1}_{\geq 0}$. For $m=2$, the same steps are applied to each $\hat{S}^{1}(j)\bm{f}$ with the scale-$j^{\prime}(>j)$ GHWT basis vectors (illustrated for $j^{\prime}=j+1$), iterating over all $j,j^{\prime}$ with $0\leq j<j^{\prime}\leq J$, yielding $S^{2}\bm{f}\in{\mathbb{R}}^{J(J+1)/2}_{\geq 0}$. This is repeated until $m$ reaches the user-specified highest order $M$ (in the figure, $M=4$ and $J=4$).
Figure 2: Schematic of the Multiscale Hodge Scattering Network for a $2$-complex with nine-elements, illustrating global pooling
@@ 3.1 Local Pooling
In general, correcting all moments up to $Q$ and of orders up to $M$ yields $Q\sum_{m=0}^{M}\binom{J+1}{m}$ features under global pooling, since the sums in (1)–(3) average over all $n$ simplices. When permutation invariance is not required (i.e., all signals are defined on a fixed complex with a known node ordering), we may omit these sums (no pooling), resulting in $Qn\sum_{m=0}^{M}\binom{J+1}{m}$ features per signal.
Between these two extremes, we can perform local pooling by averaging over regions $C^{j}_{k}$ at a chosen scale $j$, instead of averaging over the entire domain or not averaging at all. This produces a tuple of features (one per region) rather than a single value as in the global pooling case. This idea is related to windowed scattering transforms [28], but here we leverage the multiscale decomposition provided by the hierarchical partition tree, avoiding an additional window-size parameter. In local pooling, we replace the normalization factor in the average by the region size $n^{j}_{k}$ (rather than $n$), as formalized in Eq. (5). The total number of features is $Q\left(1+\sum_{m=0}^{M-1}\sum_{j=m}^{J}\binom{j}{m}K^{j}\right)=Q\left(1+\sum_{j=0}^{J}\left(\sum_{m=0}^{M-1}\binom{j}{m}\right)K^{j}\right)$,
where $K^{j}$ is the number of regions at scale $j$. Setting $K^{j}\equiv 1$ recovers the global-pooling count, whereas $K^{j}\equiv n$ recovers the no-pooling count.
We denote these transforms as $S^{m}_{j}$, where the subscript $j$ indicates the pooling scale (the scale at which the final averaging is performed). Thus, $S^{m}_{j_{\mathrm{max}}}=S^{m}$ corresponds to global pooling as in [3, 5]; $S^{m}_{0}$ denotes the transform without any pooling ($K^{0}=n$); and $S^{m}_{2}$ denotes local pooling with regions at scale $j=2$. In general,
$S^{m}_{j}\left(q,j^{(1)},\ldots,j^{(m)}\right)\bm{f}\,:=\,\Biggl\{\frac{1}{n^{j}_{k}}\sum_{i\in C^{j}_{k}}\left[\left|\Phi^{j^{(m)}}\left|\Phi^{j^{(m-1)}}\left|\cdots\left|\Phi^{j^{(1)}}\bm{f}\right|\cdots\right|\right|\right|^{q}\right]_{i}\Biggr\}_{k=0}^{K^{j}-1},$ (5)
where $n^{j}_{k}=|C^{j}_{k}|$ and $C^{j}_{k}$ is the $k$th region at scale $j$ in the partition of the input $\kappa$-simplices (i.e., partition level $p={j_{\mathrm{max}}}-j$).
Figure 3 illustrates the local-pooling version of our MHSN ($q=1$) using the same example as in Fig. 2. The key difference between this variation and the global-pooling version (i.e., averaging over all simplices) is that averages are taken over a set of disjoint local regions that together cover the entire set of simplices. As a result, instead of producing a single feature at scale $j$, we obtain one feature for each local pool at that scale. The orange triangles indicate the $2$-regions in each local pool. We construct these local pools by selecting one of the partitions of the $2$-simplices generated by the hierarchical bipartition tree used in computing the GHWT dictionary. See Appendix A and [10] for a detailed discussion of the partitioning scheme. In this figure, we use the partition defined by the support of the basis at scale j=J1=3j=J-1=3.
Figure 3: Schematic of the Multiscale Hodge Scattering Network for the same 22-complex example as in Fig. 2, illustrating local pooling